A 40.0-kg student climbs up a ladder to the roof of a building. The gravitational potential energy of the student at this point is 2.00 x 103 J. How high is the student above the ground?
1
Expert's answer
2020-12-21T11:03:03-0500

By definition, the gravitational potential energy is


"U = mgh"

where "m = 40kg" is the mass of the object, "h" is its height above the ground, and "g = 9.81m\/s^2" is the gravitational acceleration. Since "U = 2\times 10^3J", the height is


"h = \dfrac{U}{mg} = \dfrac{2\times 10^3J}{40kg\cdot 9.81m/s^2} \approx 5.1m"

Answer. 5.1 m.